Q1. Spaghetti is a family-owned restaurant in KSA. The corporate office provides 2 kinds of services (maintenance and personnel) to 3 locations (Riyadh, Jeddah and Dammam). In a recent accounting period, the cost for maintenance was $25,000 and for personnel was $15,000. Maintenance costs are allocated on the basis of square feet, and personnel costs are allocated based on the basis of number of employees. Number of Square Feet Employees Support Departments: Maintenance 1,000 15 Personnel 1,000 8 Operating Departments: Riadh 5,000 12 Jeddah 6,000 18 Dammam 3,000 10 If the restaurant uses the direct method and the costs of Riyadh location, Jeddah location and Dammam location are $15,000, $12,000 and $10,000 respectively, calculate the total costs allocated to each restaurant location.
Answer #1

ANSWER:

Total Square feet :5000 +6000 +3000 = 14000

Total Number of employees = 12+18 + 10 = 40

Support department Operating department
Maintenance Personnel Riyadh Jeddah Dammam
Direct cost 25000 15000
Allocation of Maintenance department (25000) 8929 [25000*5000/14000] 10714 [25000*6000/14000] 5357 [25000*3000/14000]
Allocation of Personnel Department (15000) 4500   [15000*12/40] 6750     [15000*18/40] 3750     [15000*10/40]
Total cost allocated 0 0 13429 17464 9107

_____________________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU**************